OpenStudy (anonymous):

PLEASE Calculus Help? (a) find the critical numbers of f (if any), (b) find the open interval(s) on which the function is increasing or decreasing, (c) apply the First Derivative Test to identify all relative extrema, and (d) use a graphing utility to confirm your results . 1. f(x)=x^3-6x^2+15 2. f(x)= x/x+1

OpenStudy (aggie12341):

1. 5^ 2. 43^4

OpenStudy (anonymous):

@aggie12341 Can you please explain?

OpenStudy (anonymous):

I don't understand the answers aggie put up?

OpenStudy (anonymous):

and honestly I don't know how to solve this. I do online schooling and don't understand half the material...

OpenStudy (anonymous):

@Loser66 could you please help?

OpenStudy (anonymous):

@satellite73

OpenStudy (loser66):

for 1/ derivative of f =?

OpenStudy (anonymous):

you mean the derivative of 1. f(x)=x^3-6x^2+15?

OpenStudy (loser66):

yes

OpenStudy (anonymous):

(x^3−6x^2+15) =(x^3)−6(x^2) =3x^2−6⋅2x =3x^2−12x

OpenStudy (anonymous):

did I do it right?

OpenStudy (loser66):

correct, now, label it, it is f'(x) = 3x^2 -12x find the roots of f'(x), I mean f'(x) =0 when x =??

OpenStudy (anonymous):

I am not sure how to find them. I tried moving the 3x^2 to the left side, dividing by 3 and so on. but I'm pretty I'm not getting the right answer

OpenStudy (loser66):

factor x out

OpenStudy (loser66):